(1)Ten more than 6 times Harold's age is the same as 4 times his age increased by 40. Find Harold's age.
(2) Using d=rt, find d for r=65mi/h and t=9hours
(3) If steve's age is decreased by 12 years, and that difference is multiplied by 3, the result is 24 years. Find steve's age

Answer by jorel1380(3719)

6H+10=4H+40
2H=30
H=15
-------------------
65 mph x 9 hours=585 miles
----------------------
3(s-12)=24
3s-36=24
3s=60
s=20
Steve is 20..